### v1

@@ 4 Computational Examples
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val SLap/SA/Gap
@@ 4.5 Meshes from Applications
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val SLap/SA/Gap
@@
From the results, it is clear that variants employing the distance Laplacian SOC matrix are generally better than those that employ $A$ as SOC matrix. The overall best method is DLap/Sgn/Val in that it successfully solves the problem in under $20$ iterations with distributed lumping over a wide range of $\theta$ values. We can also see that distributed lumping generally either reduces the iteration count or has the same iteration count as standard diagonal lumping. There are, however, two non-failure cases where distributed lumping increases the iterations. In one case, one additional iteration is taken while in another case iterations increase from $126$ to $133$. We typically find that this type of modest increase might occur for poorly converging situations. In one case, distributed lumping reduced the iteration count from $169$ to $105$ and distribution lumping requires about half as many iterations as standard lumping for the best performing combination: DLap/Sgn/Val with $\theta\geq.16$. Finally, we note that for realistic helium plume simulations depicted on the right side of Figure 11, a sophisticated control volume discretization technique is instead employed and the solution evolution requires repeated Poisson solves in a pressure-projection type of method.
@@
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val SLap/SA/Gap
@@
A dash indicates that a particular run did not converge within 500 iterations. We see that using $A$ as the SOC matrix generally leads to very poor convergence. In particular, all runs require over 250 iterations. For these poorly converging cases, we can also see a few more situations where distributed lumping might increase the number of required iterations (e.g., from 313 to 327).
The results for the jet inlet problem mirror those of the cylinder mesh, though more dramatically. In particular, it is possible to solve this problem in under $20$ iterations (e.g., with DLap/Sgn/Val and $\theta\geq.08$) while the runs that use $A$ as opposed to the distance Laplacian require hundreds of iterations. While the distributed lumping DLap/SA/Val runs are somewhat competitive, it is clearly the worst distance Laplacian based method. Similar to the 3D stretched tensor meshes, we find that DLap/SA/Gap generally performs better than DLap/Sgn/Val with standard lumping, but the opposite is true once distributed lumping is introduced. That is, distributed lumping appears to have a more positive impact on the distance Laplacian signed classical scaling method than on the cut-drop method.
@@
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val SLap/SA/Gap
@@
A dash indicates that a particular run did not converge within 500 iterations. As with previous results, using $A$ as the SOC matrix generally leads to failure. The DLap/SA/Val results with diagonal lumping are somewhat disappointing, but they improve noticeably with distributed lumping. Further, DLap/SA/Val gives by far the best operator complexities. Thus, if storage is a concern, this method combined with a low tolerance could be attractive. We note that $\theta=.01$ is indeed a fairly good choice for this problem as with $\theta=.001$ (not shown here), the iteration counts rise to 70 iterations with distributed lumping. In terms of convergence, however, the overall best method is with DLap/Sgn/Val in conjunction with distributed lumping. In particular, with $\theta\geq.16$ the number of required iterations is more than three times less than the best option with DLap/SA/Val. Once again, distributed lumping has a major impact for DLap/SA/Val and DLap/Sgn/Val, while having little influence on DLap/SA/Gap. Finally, we note that the best $\theta$ choice for DLap/Sgn/Val and DLap/SA/Gap is relatively consistent across all problems considered in this manuscript, while the best $\theta$ choice varies significantly for DLap/SA/Val, making it more problematic to define default $\theta$ values for users who do not wish to tune parameters.

### v2

@@ 4 Computational Examples
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val DLap/SA/Gap
@@ 4.5 Meshes from Applications
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val DLap/SA/Gap
@@
From the results, it is clear that variants employing the distance Laplacian SOC matrix are generally better than those that employ $A$ as SOC matrix. The overall best method is DLap/Sgn/Val in that it successfully solves the problem in under $15$ iterations with distributed lumping over a wide range of $\theta$ values. We can also see that distributed lumping generally either reduces the iteration count or has the same iteration count as standard diagonal lumping. In one case, distributed lumping reduced the iteration count from $80$ to $50$ and distribution lumping requires about half as many iterations as standard lumping for the best performing combination: DLap/Sgn/Val with $\theta\geq.16$. Finally, we note that for realistic helium plume simulations depicted on the right side of Figure 11, a sophisticated control volume discretization technique is instead employed and the solution evolution requires repeated Poisson solves in a pressure-projection type of method.
@@
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val DLap/SA/Gap
@@
We see that using the distance Laplacian as the SOC matrix generally improves the convergence rate when compared to its counterpart using the matrix $A$ as the SOC matrix. However, it should be noted that A/Sgn/Val is better than DLap/Sgn/Val when using standard diagonal lumping, but the converse is true when using distributed lumping (which is the best overall performing method).
The results for the jet inlet problem mirror those of the cylinder mesh in that the overall best method is DLap/Sgn/Val with distributed lumping, which provides converged solutions for this problem in under $20$ iterations with $\theta\geq.08$. Additionally, we see that the iteration counts are not too sensitive for the algorithm combinations that use either signed classical scaling or cut-drop (A/Sng/Val, DLap/Sng/Val or DLap/SA/Gap) when compared to the other two algorithm combinations (A/SA/Val or DLap/SA/Val) While distributed lumping DLap/SA/Val runs are somewhat competitive, it is clearly the worst distance Laplacian scheme. Similar to the 3D stretched tensor meshes, DLap/SA/Gap generally performs better than DLap/Sgn/Val with standard lumping, but the opposite is true once distributed lumping is introduced. That is, distributed lumping appears to have a more positive impact on the distance Laplacian signed classical scaling method than on the cut-drop method.
@@
$\theta$ A/SA/Val A/Sng/Val DLap/SA/Val DLap/Sng/Val DLap/SA/Gap
@@
A dash indicates that a particular run did not converge within 500 iterations. As with previous results, using $A$ as the SOC matrix generally leads to poor convergence with the noticeable exception of A/Sgn/Val which converges well for a quite large $\theta=.32$ while converging poorly for other $\theta$’s. The DLap/SA/Val results with diagonal lumping are somewhat disappointing, but they improve dramatically with distributed lumping. A/SA/Val, A/Sgn/Val, and DLap/SA/Val give by far the best operator complexities. Thus, if storage is a concern, A/Sgn/Val with a very high tolerance or DLap/SA/Val with a low tolerance and distributed lumping could be attractive. We note that $\theta=.01$ is indeed a fairly good choice for DLap/SA/Val and distributed lumping as iterations counts rise with smaller $\theta=.001$. (not shown here). In terms of convergence over a range of $\theta$’s, however, the best method is DLap/Sgn/Val in conjunction with distributed lumping. In particular, with $\theta\geq.08$ the number of required iterations is nearly three times less than the best option with DLap/SA/Val and is $2.5$ times better than A/Sgn/Val. Although the its AMG operator complexity is $1.5$ times higher than that of DLap/SA/Val, it is robust over a range of $\theta$’s and does not rely on the sign of the off-diagonal entries. Finally, we note that distributed lumping has a major impact for DLap/SA/Val and DLap/Sgn/Val, while having little influence on the other three algorithm combinations.